# Prethermalization, shadowing breakdown, and the absence of Trotterization transition in quantum circuits

### v1

@@ II.2 Kicked Ising model
There are three parameters $\tau$, $h_{\rm x}$, and $h_{\rm z}$. For generic parameters the model is nonintegrable without any local conserved quantities. We measure time in units of $\tau$, i.e., our $t$ takes integer values, and we use parameters $J=\frac{1}{2},h_{\rm x}=h_{\rm z}=1$, the same as e.g. Refs. Heyl ; Polkovnikov . Only in section III.4.2 on transport will we exclusively use different ones in order to compare with values of diffusion constant found in literature.
@@ II.3 Spectrum of the truncated propagator
At $\tau>2$ the spectrum qualitatively changes and is made out of 3 rings. Point $\tau=\pi$ is special because it is integrable (two otherwise different integrable points merge into one because of $h_{\rm x}=h_{\rm z}=1$RP24 ; prosen07 : singular values of $M(k)$ are all $1$ ($U_{\rm x}=\prod_{j}(-{\rm i}\sigma^{\rm x}_{j}),U_{\rm z}=\prod_{j}(-{\rm i}\sigma^{\rm z}_{j})\prod_{p}(1-{\rm i}\sigma^{\rm z}_{p}\sigma^{\rm z}_{p+1})/\sqrt{2}$), while eigenvalues of $M(k)$ are $10\cdot 4^{r-3}$ equal to $+1$, the same number is $-1$, while the rest is 00 (i.e., a fraction $5/24$ of all is $+1$). The fact that nonzero eigenvalues for $\tau\to\pi$ bunch around ±$\pm 1$ is a signature of a discrete time crystal vedika16 ; else16 ; curt16 , specifically of the model discussed in e.g. Ref. se21 . As we shall see, expectation values of some observables will exhibit oscillations with period $2$ (Sec.III.5). Spectrum of $M(k)$ will therefore tell us a lot of interesting things.
@@ III.1 Leading Ruelle-Pollicott resonance and prethermalization
The most important quantity is the eigenvalue of $M(k)$ with the largest modulus. It determines the long-time decay rate of correlation functions of generic observables (those not orthogonal to the corresponding eigenvector) under evolution by circuit $U_{\tau}$. Furthermore, as we shall se, it also determines the duration of the prethermalization phase, as well as a range of times and length-scales on which one can observe transient diffusion during the prethermalization phase, being a Floquet remnant of energy diffusion in the Hamiltonian $H_{0}$. Other methods have been used to find “slowest” observables, like minimizing the commutator with $U_{\tau}$ mari15 (fastest decay at short times), see also Refs. motrunich17 ; izotova23 , or using time averaging marcin15 .
@@
The exact Floquet Hamiltonian $H_{\rm F}$ is in general nonlocal and can not be obtained by e.g. BCH expansion because the series is divergent, see Appendix B. However, upto prethermalization time $T$ we can write the propagator approximately in terms of an almost conserved operator $H_{\rm F}^{\prime}$. While one can in principle obtain $H_{\rm F}^{\prime}$ via a BCH expansion, such a procedure is fraught with a certain freedom of what is the “optimal” truncation of an otherwise divergent series. The truncated operator propagator provides a more direct way – $H_{\rm F}^{\prime}$ is nothing but the eigenvector corresponding to $\lambda_{1}$.
@@
Important for our discussion of shadowing is behavior after prethermalization ends, $t\gg T$. We can see in Fig. 2D that at finite $L$ a 2nd lower plateau is reached, which however decreases to 00 as on increases $L$. In the thermodynamic limit, where one has to first send $L\to\infty$ and only then $t\to\infty$, there is therefore no plateau in correlation functions, nor in the magnetization itself. This is different than what has been observed in Refs. Heyl ; Polkovnikov ; Bukov – different conclusions there were a consequence of using a finite averaging time over which one calculated e.g. magnetization (i.e., the averaging time was smaller than the prethermalization time $T$ (15), or, using too small $L$ for the chosen averaging time). In the TDL of chaotic systems there is therefore no Trotterization transition due to mysterious “quantum localization”. What was observed in Ref. Heyl is a simple manifestation of a single almost conserved operator due to prethermalization. Let us remark that the Trotterization transition in integrable systems vernier23 , for instance in XXZ-type integrable circuits, is of a completely different origin – it is a consequence of different phases between which the whole analytical integrability structure changes and thereby also physical properties kpzwall .
@@
Figure 4: Prethermalization vs. Heisenberg time. Finite-size asymptotic value of magnetization correlation function (e.g., from Fig. 2C we can read $C_{Z}(\infty)\approx 0.1$ for $L=24$ and $\tau=1.0$, reached at $t>10^{5}$) is shown for various sizes $L=14-28$ and $\tau$. To observe the correct thermodynamic limit $\lim_{t\to\infty}\lim_{L\to\infty}$, in which $C_{Z}(\infty)$ decays as $\propto 1/2^{L}$ (green line), one must have $T\ll t_{\rm H}$ – the prethermalization time $T=1/\Delta$ has to be smaller than the Heisenberg time $t_{\rm H}\sim 2^{L}$.
@@ III.2 Finite-size effects and the Heisenberg time
Being careful about $t_{\rm H}$ is especially important in systems with long time scales, for instance, in our case the prethermalization time $T$ is exponentially large in $1/\tau$ (e.g., Fig. 2C), another case is systems with disorder where dynamics gets exponentially slow with increasing disorder strength panda19 . For instance, in Fig. 2D we can see that at $L=16$ it indeed seems that even at $t\sim 10^{6}$ correlations do not decay, and one would be tempted to conclude that one is in a non-mixing regime. However, that is not correct. For larger $L$ correlations do start to decay. The reason for non-decay at $L=16$ is simply that one is not in the regime $T\ll t_{\rm H}$. In order to see the correct physics obtained in the thermodynamic limit one should have
@@
We numerically demonstrate correctness of Eq.(18) in Fig. 4. From numerical simulation data we read out the value of $C_{Z}(t)$ at large time, where a 2nd plateau is reached for finite $L$. Plotting this value $C_{Z}(t\to\infty)$ as a function of scaled time $t_{\rm H}/T$, where $T$ is read of from Fig. 2C, we get a good collapse of data for all $L$ and $\tau$ (i.e., $T$). We indeed see that in order to really see decaying correlations, where $C_{Z}(\infty)\sim 1/2^{L}$, one must have $t_{\rm H}\gg T$. Different conclusions in Refs. Heyl ; Polkovnikov ; Bukov are due to violation of this condition. Regarding the stability of expectation of $H_{0}$, the good news is that the prethermalization time is exponentially large in $1/\tau$ and so one does not need excessively small time step $\tau$ to observe long stability, however, if one wants to study the TDL one should still respect $T\ll t_{\rm H}$ and therefore also not take a too small $\tau$ at a given $L$; the minimal $\tau$ scales as $1/L$ (18).
@@ III.3 Fast decay and no prethermalization of generic observables
All other observables that are orthogonal to $H_{\rm F}^{\prime}$ will deteriorate on timescales that are inline with Trotter errors, that is, much faster than the prethermalization time $T$. Note that while a typical observable from the even zero momentum sector will have a nonzero overlap with $H_{\rm F}^{\prime}$ (which though, depending on what we mean by typical, i.e. the measure, can be small), any observable with a nonzero $k$ will have zero overlap. Therefore, allowing observables with nonzero momentum a typical observable will not exhibit prethermalization. This is numerically demonstrated in Fig. 5B. We pick a staggered magnetization $S\equiv\sum_{j}(-1)^{j}\sigma^{\rm z}_{j}$ and observe that the correlation function for finite $\tau$ starts to differ from the one obtained for Hamiltonian evolution generated by $H_{0}$ much more rapidly than e.g. $H_{0}$ or $Z$ in Fig. 2D which do exhibit prethermalization. For the largest $\tau=0.8$ shown, for which the staggered magnetization visibly deviates already at $t\approx 22$, the total magnetization $Z$ would follow the Hamiltonian result upto very large prethermalization $T\approx 10^{8}$. Typical observable is therefore much more sensitite to Trotterization errors because it does not exhibit prethermalization.
@@ III.4.1 Finite-times
Upto large time $T$ the operator $H_{\rm F}^{\prime}$ behaves as if it would be conserved. Similarly as one can study energy transport under Hamiltonian evolution by $H_{0}$, one can also ask about transport properties of $H_{\rm F}^{\prime}$ in the circuit $U_{\tau}$ for $t\ll T$ when $H_{\rm F}^{\prime}$ is quasi conserved.
@@ III.4.2 Hamiltonian limit
Figure 7: Diffusion constant for Hamiltonian tilted Ising model. (A) Using the truncated propagator in the limit of small $\tau$ and $k$, getting $D=\lambda_{1}(k)/(k^{2}\tau)$ at $k=0.01$. (B) Lindblad NESS with $L=100$ and high-temperature boundary parameters. Shown is the energy density profile (the NESS current $j_{p}$ is independent of site $p$). (C) Green-Kubo and the energy current autocorrelation function. All three methods give consistently the same value of $D$ (${\cal J}=1.0,h_{\rm x}=1.4,h_{\rm z}=0.9045$).
For small $\tau$ the prefactor in front of $\sim k^{2}$ is trivially proportional to $\tau$ because in the Hamiltonian limit of $\tau\to 0$ one should recover the diffusion constant of the Hamiltonian $H_{0}$ which is measured in units of physical time $t\tau$, and not in time steps $t$ as obtained from the gap of $U_{\tau}$. To check that $D$ obtained in such a way from $M(k)$ is indeed equal to the diffusion constant of $H_{0}$ we have calculated $D$ directly for the Hamiltonian by two different methods, one is the Lindblad equation with boundary driving where one studies a true nonequilibirum steady states (NESS), while the other is the Green-Kubo formula using the equilibrium current autocorrelation function.
In order to facilitate comparison with existing data we here take slightly different parameters, namely, we use $H_{0}^{\prime}=\sum_{j}{\cal J}\sigma^{\rm z}_{j}\sigma^{\rm z}_{j+1}+h_{\rm z}\sigma^{\rm z}_{j}+h_{\rm x}\sigma^{\rm x}_{j}$, with ${\cal J}=1$ and $h_{\rm x}=1.4,h_{\rm z}=0.9045$. A number of recent papers tibor22 ; bard24 ; robin24 ; white24 has values of $D$ at infinite temperature for those parameters, with values in the range $D=1.38-1.55$, with the most accurate seemingly in Ref. white24 at $D\approx 1.445$. Our results for $D$ obtained from a small $k$ behavior of $\lambda_{1}(k)$ of the truncated propagator are in Fig. 7A. Taking small $\tau=0.05$ we get $D=\lambda_{1}(k=0.01)/(k^{2}\tau)\approx 1.462$ for truncation size $r=12$. We can also see that at $r=12$ there is still small decay with $r$ so even larger $r$ would be needed to obtain complete convergence (numerically this is expensive because the gap is very small, of order $10^{-5}$ at $\tau=0.05$, and the size of $M$ is already $\approx 10^{7}$ at $r=12$) so we extrapolate last few points and estimate that the true value is $D=1.44-1.46$ (we give it a rather generous error bar because the extrapolation is not super nice).
@@ III.5 Prethermalization at large $\tau$ and DTC
Numerically calculating the correlation function, for instance of $H_{0}$ (10, we indeed see (Fig. 8) long prethermalization plateau, exactly as was the case for small $\tau$. The gap again perfectly predicts the thermalization time $T=1/\Delta$, with the scaling being (Fig. 2C) $T\approx 10^{-8}\exp{(21/|\pi-\tau|)}$. This case of large $\tau$ is also interesting for another reason, namely, because the propagator is close to a root-of-identity we can expect an almost periodic behavior in time, that is, breaking of the time-translation symmetry, i.e., a discrete time crystal. It has been known Nayak that prethermalization is a way to obtain DTC in homogeneous systems, however not in (short-ranged) 1-dimensional systems. In our case we can observe that the total magnetization oscillates (Fig. 8) with a period $2$. Similar behavior has been observed before in the kicked Ising model, see Refs. marin21 ; jaksch19 ; pizzi20 ; se21 , seemingly at odds with a 1-dimensional nature of the system (for classical 1-d DTC see Ref. li25 ). We in fact have three different regimes in the behavior of $Z(t)$ (see Fig. 8): (i) initial fast decay of magnetization until a minimum at $t\approx L$ (see Ref. deLuca for a detailed discussion of this regime), (ii) intermediate prethermal times $L\ll t\ll T$ where one has a DTC-like behavior, though the size of the plateau in the average spin magnetization $Z(t)/L$ here decays with $L$ as $\sim 1/L$, and (iii) asymptotic times $t\gg T$ where prethermalization breaks down and $Z(t)$ decays to a value that is exponentially small in $L$. Although one does not have a true DTC in the thermodynamic limit in regime (ii), it could still be of (experimental) interest as the plateau for specific initial states, like ${|1\ldots 1\rangle}$, is still relatively large, say $>0.1$ at $L=32$, despite being rather far away from the exact DTC at $\tau=\pi$. Previous works marin21 ; jaksch19 ; pizzi20 ; se21 mostly focused on smaller deviations from $\pi$, where one has to be extra careful about finite size effects, see also discussion in Ref. deLuca . Because the truncated propagator gives an easy way to obtain the prethermalization duration $T=1/\Delta$ it also gives an exact prediction on the robustness of potential DTCs, avoiding confusing a finite-$L$ DTC with an exact one in the TDL. As discussed (Sec. III.2), if one is interested in the TDL one should be aware of $T$ and of the Heisenberg time, and look at appropriate $L$ and $t$.

### v2

@@ II.2 Kicked Ising model
There are three parameters $\tau$, $h_{\rm x}$, and $h_{\rm z}$. For generic parameters the model is nonintegrable without any local conserved quantities. We measure time in units of $\tau$, i.e., our $t$ takes integer values, and we use parameters $J=\frac{1}{2},h_{\rm x}=h_{\rm z}=1$, the same as e.g. Refs. Heyl ; Polkovnikov . Only in section III.4.2 on transport will we use different ones in order to compare with values of the diffusion constant found in literature.
@@ II.3 Spectrum of the truncated propagator
At $\tau>2$ the spectrum qualitatively changes and is made out of 3 rings. Point $\tau=\pi$ is special because it is integrable (two otherwise different integrable points merge into one because of $h_{\rm x}=h_{\rm z}=1$RP24 ; prosen07 : singular values of $M(k)$ are all 11 ($U_{\rm x}=\prod_{j}(-{\rm i}\sigma^{\rm x}_{j}),U_{\rm z}=\prod_{j}(-{\rm i}\sigma^{\rm z}_{j})\prod_{p}(1-{\rm i}\sigma^{\rm z}_{p}\sigma^{\rm z}_{p+1})/\sqrt{2}$), while eigenvalues of $M(k)$ are $10\cdot 4^{r-3}$ equal to +1+1, the same number is 1-1, while the rest is 0 (i.e., a fraction $5/24$ of all is +1+1). The fact that nonzero eigenvalues for $\tau\to\pi$ bunch around $\pm 1$ is a signature of a discrete time crystal vedika16 ; else16 ; curt16 , specifically of the model discussed in e.g. Ref. se21 . As we shall see, expectation values of some observables will exhibit oscillations with period $2$ (Sec.III.5). Spectrum of $M(k)$ will therefore tell us a lot of interesting things.
@@ III.1 Leading Ruelle-Pollicott resonance and prethermalization
The most important quantity is the eigenvalue of $M(k)$ with the largest modulus. It determines the long-time decay rate of correlation functions of generic observables (those not orthogonal to the corresponding eigenvector) under evolution by circuit $U_{\tau}$. Furthermore, as we shall see, it also determines the duration of the prethermalization phase, as well as a range of times and length-scales on which one can observe transient diffusion during the prethermalization phase, being a Floquet remnant of energy diffusion in the Hamiltonian $H_{0}$. Other methods have been used to find “slowest” observables, like minimizing the commutator with $U_{\tau}$ mari15 (fastest decay at short times), see also Refs. motrunich17 ; izotova23 , or using time averaging marcin15 .
@@
The exact Floquet Hamiltonian $H_{\rm F}$ is in general nonlocal and can not be obtained by e.g. BCH expansion because the series is divergent, see Appendix B. However, up-to prethermalization time $T$ we can write the propagator approximately in terms of an almost conserved operator $H_{\rm F}^{\prime}$. While one can in principle obtain $H_{\rm F}^{\prime}$ via a BCH expansion, such a procedure is fraught with a certain freedom of what is the “optimal” truncation of an otherwise divergent series. The truncated operator propagator provides a more direct way – $H_{\rm F}^{\prime}$ is nothing but the eigenvector corresponding to $\lambda_{1}$.
@@
Important for our discussion of shadowing is behavior after prethermalization ends, $t\gg T$. We can see in Fig. 2D that at finite $L$ a 2nd lower plateau is reached, which however decreases to 0 as on increases $L$. In the thermodynamic limit, where one has to first send $L\to\infty$ and only then $t\to\infty$, there is therefore no plateau in correlation functions, nor in the magnetization itself. This is different than what has been observed in Refs. Heyl ; Polkovnikov ; Bukov – different conclusions there were a consequence of using a finite averaging time over which one calculated e.g. magnetization (i.e., the averaging time was smaller than the prethermalization time $T$ (15), or, using too small $L$ for the chosen averaging time). In the TDL of chaotic systems there is therefore no Trotterization transition due to mysterious “quantum localization”. What was observed in Ref. Heyl is a simple manifestation of a single almost conserved operator due to prethermalization. Let us remark that the Trotterization transition in integrable systems vernier23 , for instance in XXZ-type integrable circuits, is of a completely different origin – it is a consequence of different phases between which the whole analytical integrability structure changes and thereby also physical properties kpzwall .
@@
Figure 4: Prethermalization vs. Heisenberg time. Finite-size asymptotic value of magnetization correlation function (e.g., from Fig. 2C we can read $C_{Z}(\infty)\approx 0.1$ for $L=24$ and $\tau=1.0$, reached at $t>10^{5}$) is shown for various sizes L=1428L=14-28 and $\tau$. To observe the correct thermodynamic limit $\lim_{t\to\infty}\lim_{L\to\infty}$, in which $C_{Z}(\infty)$ decays as $\propto 1/2^{L}$ (green line), one must have $T\ll t_{\rm H}$ – the prethermalization time $T=1/\Delta$ has to be smaller than the Heisenberg time $t_{\rm H}\sim 2^{L}$.
@@ III.2 Finite-size effects and the Heisenberg time
Being careful about $t_{\rm H}$ is especially important in systems with long time scales, for instance, in our case the prethermalization time $T$ is exponentially large in $1/\tau$ (e.g., Fig. 2C), another case is systems with disorder where dynamics gets exponentially slow with increasing disorder strength panda19 . For instance, in Fig. 2D we can see that at $L=16$ it indeed seems that even at $t\sim 10^{6}$ correlations do not decay, and one would be tempted to conclude that one is in a non-mixing regime. However, that is not correct. For larger $L$ correlations do start to decay. The reason for non-decay at $L=16$ is simply that one is not in the regime $T\ll t_{\rm H}$. This is another example where relying on an exact diagonalization, either full spectrum or just an often-used diagonal approximation, would completely incorrectly diagnose the long-time dynamics in the thermodynamic limit (e.g., plateau in $L=16$ data in Fig. 2D).
In order to see the correct physics in the thermodynamic limit one should have
@@
We numerically demonstrate correctness of Eq.(18) in Fig. 4. From numerical simulation data we read out the value of $C_{Z}(t)$ at large time, where a 2nd plateau is reached for finite $L$. Plotting this value $C_{Z}(t\to\infty)$ as a function of scaled time $t_{\rm H}/T$, where $T$ is read of from Fig. 2C, we get a good collapse of data for all $L$ and $\tau$ (i.e., $T$). We indeed see that in order to really see decaying correlations, where $C_{Z}(\infty)\sim 1/2^{L}$, one must have $t_{\rm H}\gg T$. Different conclusions in Refs. Heyl ; Polkovnikov ; Bukov are due to violation of this condition. Regarding the stability of expectation of $H_{0}$, the good news is that the prethermalization time is exponentially large in $1/\tau$ and so one does not need excessively small time step $\tau$ to observe long stability, however, if one wants to study the TDL one should still respect $T\ll t_{\rm H}$ and therefore also not take a too small $\tau$ at a given $L$; the minimal meaningful timestep $\tau$ scales in a finite system as $1/L$ (18).
@@ III.3 Fast decay and no prethermalization of generic observables
All other observables that are orthogonal to $H_{\rm F}^{\prime}$ will deteriorate on timescales that are inline with Trotter errors, that is, much faster than the prethermalization time $T$. Note that while a typical observable from the even zero momentum sector will have a nonzero overlap with $H_{\rm F}^{\prime}$ (which though, depending on what we mean by typical, i.e. the measure, can be small), any observable with a nonzero $k$ will have zero overlap. Therefore, allowing observables with nonzero momentum a typical observable will not exhibit prethermalization. This is numerically demonstrated in Fig. 5B. We pick a staggered magnetization $S\equiv\sum_{j}(-1)^{j}\sigma^{\rm z}_{j}$ and observe that the correlation function for finite $\tau$ starts to differ from the one obtained for Hamiltonian evolution generated by $H_{0}$ much more rapidly than e.g. $H_{0}$ or $Z$ in Fig. 2D which do exhibit prethermalization. For the largest $\tau=0.8$ shown, for which the staggered magnetization visibly deviates already at $t\approx 22$, the total magnetization $Z$ would follow the Hamiltonian result up-to very large prethermalization $T\approx 10^{8}$. Typical observable is therefore much more sensitive to Trotterization errors because it does not exhibit prethermalization.
@@ III.4.1 Finite-times
Up-to large time $T$ the operator $H_{\rm F}^{\prime}$ behaves as if it would be conserved. Similarly as one can study energy transport under Hamiltonian evolution by $H_{0}$, one can also ask about transport properties of $H_{\rm F}^{\prime}$ in the circuit $U_{\tau}$ for $t\ll T$ when $H_{\rm F}^{\prime}$ is quasi conserved.
@@
Figure 7: Diffusion constant for Hamiltonian tilted Ising model. (A) Using the truncated propagator in the limit of small $\tau$ and $k$, getting $D=[1-\lambda_{1}(k)]/(k^{2}\tau)$ at $k=0.02$. (B) Lindblad NESS with $L=100$ and high-temperature boundary parameters. Shown is the energy density profile (the NESS current $j_{p}$ is independent of site $p$). (C) Green-Kubo and the energy current autocorrelation function. All three methods give consistently the same value of $D$ (${\cal J}=1.0,h_{\rm x}=1.4,h_{\rm z}=0.9045$).
@@ III.4.2 Hamiltonian limit
For small $\tau$ the prefactor in front of $\sim k^{2}$ is trivially proportional to $\tau$ because in the Hamiltonian limit of $\tau\to 0$ one should recover the diffusion constant of the Hamiltonian $H_{0}$ which is measured in units of physical time $t\tau$, and not in time steps $t$ as obtained from the gap of $U_{\tau}$. To check that $D$ obtained in such a way from $M(k)$ is indeed equal to the diffusion constant of $H_{0}$ we have calculated $D$ directly for the Hamiltonian by two other methods, one is the Lindblad equation with boundary driving where one studies a true nonequilibrium steady states (NESS), while the other is the Green-Kubo formula using the equilibrium current autocorrelation function.
In order to facilitate comparison with existing data we here take slightly different parameters, namely, we use $H_{0}^{\prime}=\sum_{j}{\cal J}\sigma^{\rm z}_{j}\sigma^{\rm z}_{j+1}+h_{\rm z}\sigma^{\rm z}_{j}+h_{\rm x}\sigma^{\rm x}_{j}$, with ${\cal J}=1$ and $h_{\rm x}=1.4,h_{\rm z}=0.9045$. A number of recent papers has values of $D$ at infinite temperature for those parameters, with values in the range D=1.381.55D=1.38-1.55, specifically, Ref. tibor22 gets $D\approx 1.40$, Ref. bard24 $D\approx 1.55$, Krylov method of Ref. robin24 $D\approx 1.38$, with the most accurate value seemingly in Ref. white24 at $D\approx 1.445$. Our results for $D$ obtained from a small $k$ behavior of $\lambda_{1}(k)$ of the truncated propagator are in Fig. 7A. Taking small $\tau=0.05$ we get $D=(1-\lambda_{1}(k=0.02))/(k^{2}\tau)\approx 1.454$ for truncation size $r=13$, while doing extrapolation footconv to $r\to\infty$ gets us $D\approx 1.443\pm 0.005$ (similar value is obtained for $k=0.01$). We see that our estimate agrees well with Ref. white24 , while other methods are less accurate. For larger $\tau=0.1$, when the Trotterization is not negligible, the converged value is a bit larger and reflects a finite-time diffusion constant in a circuit with long prethermalization time.
@@ III.5 Prethermalization at large $\tau$ and DTC
Numerically calculating the correlation function, for instance of $H_{0}$ (10, we indeed see (Fig. 18) long prethermalization plateau, exactly as was the case for small $\tau$. The gap again perfectly predicts the thermalization time $T=1/\Delta$, with the scaling being (Fig. 2C) $T\approx 10^{-8}\exp{(21/|\pi-\tau|)}$. This case of large $\tau$ is also interesting for another reason, namely, because the propagator is close to a root-of-identity we can expect an almost periodic behavior in time, that is, breaking of the time-translation symmetry, i.e., a discrete time crystal. It has been known Nayak that prethermalization is a way to obtain DTC in homogeneous systems, however not in (short-ranged) 1-dimensional systems. In our case we can observe that the total magnetization oscillates (Fig. 18) with a period $2$. Similar behavior has been observed before in the kicked Ising model, see Refs. marin21 ; jaksch19 ; pizzi20 ; se21 , seemingly at odds with a 1-dimensional nature of the system (for classical 1-d DTC see Ref. li25 ). We in fact have three different regimes in the behavior of $Z(t)$ (see Fig. 18): (i) initial fast decay of magnetization until a minimum at $t\approx L$ (see Ref. deLuca for a detailed discussion of this regime), (ii) intermediate prethermal times $L\ll t\ll T$ where one has a DTC-like behavior, though the size of the plateau in the average spin magnetization $Z(t)/L$ here decays with $L$ as $\sim 1/L$, and (iii) asymptotic times $t\gg T$ where prethermalization breaks down and $Z(t)$ decays to a value that is exponentially small in $L$. Although one does not have a true DTC in the thermodynamic limit in regime (ii), it could still be of (experimental) interest as the plateau for specific initial states, like $|1\ldots 1\rangle$}, is still relatively large, say >0.1>0.1 at $L=32$, despite being rather far away from the exact DTC at $\tau=\pi$. Previous works marin21 ; jaksch19 ; pizzi20 ; se21 mostly focused on smaller deviations from $\pi$, where one has to be extra careful about finite size effects, see also discussion in Ref. deLuca . Because the truncated propagator gives an easy way to obtain the prethermalization duration $T=1/\Delta$ it also gives an exact prediction on the robustness of potential DTCs, avoiding confusing a finite-$L$ DTC with an exact one in the TDL. As discussed (Sec. III.2), if one is interested in the TDL one should be aware of $T$ and of the Heisenberg time, and look at appropriate $L$ and $t$.
Figure 9: 2D kicked Ising model on a square lattice (illustrated in (B)). (A) dependence of the gap on $\tau$. Three crosses mark the empirical decay used as dashed black exponential curves for three $\tau$ in (D), while brown squares are best fitted exponentials to numerical data for largest size $L=6\times 5$ in (D). (C) shows spectra of $M(\mathbf{k}=0)$ for truncation $r=3\times 2$.
@@
III.6 2D kicked Ising model
So far we have demonstrated the ideas on the 1D kicked Ising model. The methods and conclusions though are model independent and here we briefly comment and show data for some other models.
One could specifically wonder if the kicked Ising model is generic enough, for instance, it has a single layer of commuting 2-qubit gates (Fig. 3), whereas one might be interested in say a brickwall circuit where one has 2 layers of 2-qubit gates (with gates commuting within each layer). To that end we have implemented a model where on top of kicked Ising gates we add another layer of commuting $xx$ gates, thereby making the model more generic. In the limit of small $\tau$ such a model corresponds to a 1D XX model in a tilted magnetic field. More detailed results are presented in Appendix C, here we just summarize main findings. At small $\tau$ one again finds prethermalization with its timescale given by the leading Ruelle-Pollicott resonance of the truncated propagator. Close to $\tau=\pi$ things though are different than in the kicked Ising model. While the timescale again diverges, it does so only quadratically in $1/(\tau-\pi)$ (see Fig. 16D), and not exponentially like in the kicked Ising model. This is so because the kicked XX model is not a DTC at $\tau=\pi$. In fact, behavior is similar to the kicked Ising model at small $\tau$ in sectors with nonzero quasimomentum $k$ where there are no almost-conserved operators (Fig. 5). From the momentum dependence of $\lambda_{1}$ one can again get the energy diffusion constant which agrees with the one obtained from the integral of the current autocorrelation function via the Green-Kubo formula (Fig. 17).
The truncated operator propagator can be used also in 2D models and we focus here on an example of a 2D kicked Ising model. Spins are now located on a 2D square lattice, with a one-step propagator being
$U_{\tau}$ $=$ $U^{\prime}_{\rm zz}U^{\prime}_{\rm z}U^{\prime}_{\rm x},$ (23)
$U^{\prime}_{\rm x}$ $=$ $\prod_{i,j}{\rm e}^{-{\rm i}\tau h_{\rm x}\sigma^{\rm x}_{i,j}},$
$U^{\prime}_{\rm z}$ $=$ $\prod_{i,j}{\rm e}^{-{\rm i}\tau h_{\rm z}\sigma^{\rm z}_{i,j}},$
$U^{\prime}_{\rm zz}$ $=$ ${\rm e}^{-{\rm i}\tau\sum_{\rm n.n.}\sigma^{\rm z}_{i,j}\sigma^{\rm z}_{i^{\prime},j^{\prime}}},$
where the sum in the $U^{\prime}_{\rm zz}$ is over nearest-neighbors on a 2D square lattice (Fig. 9B). Similar to the 1D model we calculate the operator propagator in an infinite system truncated to a basis of local operators supported on $r$ sites. Because we are in 2D, the local basis operators are now from $r$ sites on an $n\times m=r$ grid. We show results for the largest eigenvalue for the $r=2\times 2$ and $3\times 3$ truncations in Fig. 9(A). Because one has translational invariance in $x$ as well as in $y$ direction, resulting in two conserved quasimomentum components $k_{x}$ and $k_{y}$, construction of independent basis elements is a bit more involved. One has to ensure that all basis elements are independent, i.e., that they can not be obtained from each other by translations. In 1D it was enough to fix e.g. the first-site operator to be a non-identity, in 2D this is not enough. Details can be found in Appendix D, let us just mention that the size of the basis approaches $4^{r=mn}$ for large $m$ and $n$. We also use in-situ numerical propagation of local basis operators so that we can easily study $r=3\times 3$ truncation.
Dynamics of the 2D model probably warrants a separate study, here we focus on the main things in the zero quasimomentum sector ($k_{x}=k_{y}=0$), setting also $h_{\rm z}=h_{\rm z}=1$ and varying only $\tau$ (note that compared to the 1D kicked Ising model we do not use factors $1/2$ or $1/4$ in the exponents). At small $\tau$ one finds prethermalization with the prethermalization time being exponentially large in $1/\tau$, finding no Trotterization transition (Fig. 9A). The decay time-scale at larger $\tau=0.60$ obtained from the exact numerical simulation of finite systems of size up-to $L=6\times 5$ spins with periodic boundary conditions agrees with the one obtained from the truncated propagator at truncation size $r=3\times 3$ (Figs. 9A,D). At smaller $\tau$ one can see a rapid increase of the prethermalization time, e.g. between $\tau=0.45$ and $\tau=0.35$ it increases by almost $2$ decades, though small linear truncation size ($2$ or $3$) is not yet large enough to get a precise value of the prethermalization time (this is actually similar as in 1D). Nevertheless, one can see a clear rapid decrease of the gap with decreasing $\tau$ (Fig. 9A). Similarly as in 1D relying on small systems, e.g. with $L=4\times 4=16$ spins, would incorrectly indicate a plateau at large times (green curves in Fig. 9D).
Figure 10: Transition in the behavior of magnetization correlation function in the 2D kicked Ising model around $\tau\approx 1.3$. Top set of data (points) is for $\tau=1.35$, while the lower ones (curves) are for $\tau=1.25$.
There are two other interesting points worth highlighting. The first is that at $\tau=\pi/4$ all eigenvalues of $M(\mathbf{k}=0)$ are zero, i.e., the gap shown in Fig. 9A is exactly zero. While this is not specific to 2D, and has been already noted in the 1D kicked Ising model RP24 , it means that all autocorrelation functions of extensive local operators decay immediately. We state some further properties of $M$ at $\tau=\pi/4$ and $\tau=\pi/2$ in Appendix D.
The second point is an apparent transition in the dynamics around $\tau\approx 1.3$. Namely, for $\tau>\pi/4$ one notices that the largest absolute eigenvalue becomes negative, as opposed to being positive for $\tau<\pi/4$ (see e.g. the spectrum for $\tau=1.0$ in Fig. 9C). As a consequence correlation functions will oscillate between negative and positive values between even and odd times. What is even more interesting is that around $\tau\approx 1.3$ there seems to be a qualitative change in the spectrum (Fig. 9C) which does not anymore have a noisy ring or disk core (like in both 1D models). This change is reflected also in the decay of correlation functions. In Fig. 10 we can see that at $\tau=1.25$ the autocorrelation function of magnetization has two distinct decay regimes: up-to $t\approx 50$ its decay fast in an oscillatory manner ($C_{\rm Z}$ at even and odd times have opposite signs), whereas for $t>50$ it is positive and decays slower – perhaps still exponentially, though our largest system size is not large enough to verify that with confidence. On the other hand, for $\tau=1.35$ the decay is much slower, has a single regime ($C_{\rm Z}$ is always positive), with the asymptotic decay perhaps being more likely a power law rather than exponential. This could be due to a continuum of eigenvalues approaching $1$ in modulus (see Ref. RP24 for an example where a continuum of eigenvalues (a branch cut) causes a power law decay).